(12-33) Grannys pays a tax rate of 25%. Granny is considering the purchase of a new Turbo Churn for $25,000. The churn is ause excel and show functions
Answer #1

rate positively ..

We have to compute the NPV of the project
Annual depreciation = (25000-5000)/4 5000
year 0 1 2 3 4
A Initial investment -25000
i Saving 8000 8000 8000 8000
ii Depreciation 5000 5000 5000 5000
iii=i-ii Profit before tax 3000 3000 3000 3000
iv=iii*25% Tax @ 25% 750 750 750 750
v=iii-iv Profit after tax 2250 2250 2250 2250
B=v+ii Cash flow 7250 7250 7250 7250
C Salvage value after tax = 5000
D=A+B+C Net cash flow -25000 7250 7250 7250 12250
E PVIF @ 10% 1 0.909091 0.826446 0.751315 0.683013
F=D*E Present value= (25,000.00) 6,590.91 5,991.74 5,447.03 8,366.91 1,396.59
Therefore, NPV =      1,396.59